Shift report 20000809 00:00-08:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Wed Aug 09 2000 - 08:11:14 EDT

  • Next message: Peter H. L. Christiansen: "Re: Shift report 20000809 16:00-24:00"

    ------------------------------------------------------------
    Date: Wed Aug  9 08:11:14 2000, Shift: 00:00-08:00
    Supervisor: DO, with: ZIS, DB
    Log book#4 pages 38 - 40
    
    Shift Summary:
    ------------------------------------------------------------
    1. stable beam during all night, MC stopped it around 6:20
       fire alarm in RHIC tunnel. Preparing for reinjection 
       of beam. 7:40 => new fire alarm
    2. Run 2145 (started in previous shift) and a few test runs
    3. Around 1:50, noticed event rate = 0 in DAQ. Stopped the
       run. Got a lot of EventBuilder error. MC called to ask 
       why scalers were stuck. Reset the scalers in LumiScaler.
       Individual ZDC and BB were counting, the others remained
       at 0. Visual scalers in FEH were stuck too. Reset them 
       but remained at 0. Called Konstantin and Dana. 
       Dana came around 3:00. Cycled power of some crates NIM 
       bins and got scalers back. See logbook for details. 
       Got the DAQ work when only camac and fb4 partitions were
       selected => TPC and DC partitions screw the DAQ. 
       Sent an email to Bob Scheetz to check TPC readout.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Wed Aug 09 2000 - 08:14:30 EDT